(4a^3b^3)(4ab^2)^2
Answer by RAY100(1637)

(4a^3b^3)(4ab^2)^2
.
(4a^3b^3)(16a^2b^4),,,,,(b^2)^c= b^2c
.
64 a^5b^7,,,,,,,a^b*a^c=a^(b+c)
.
.
check let a=2 b=3,,both origin and answer = 4,478,976,,,,ok